OpenStudy (anonymous):

Find the polynomial function with roots 1, –2, and 5

OpenStudy (mertsj):

Could you solve this equation: (x-3)(x+2)(x-6)=0?

OpenStudy (mertsj):

@Lesha ????

OpenStudy (anonymous):

yes i could (: x=3, x=-2 and x=6. right ?

OpenStudy (mertsj):

Very good.

OpenStudy (mertsj):

Now could you reverse the process? If I said, x = 3, x = -2 and x = 6, could you write the factors?

OpenStudy (anonymous):

im not sure what that means ?

OpenStudy (mertsj):

I said: (x-3)(x+2)(x-6)=0 You correctly said that x=3, x=-2 and x = 6

OpenStudy (mertsj):

So let's start at the end and go backwards: x = 3, x = -2 and x = 6 so the equation must have been: (x-3)(x+2)(x-6)=0

OpenStudy (anonymous):

ok ii get that part.

OpenStudy (anonymous):

now what ?

OpenStudy (mertsj):

So according to your problem: x=1, x = -2 and x = 5 So the equation must have been what?